# Gender differences in collaboration and career progression in physics

### v1

@@ 5.2 Duration of career stages
We examine the impact of collaborative behaviour and gender on career progression using accelerated failure time (AFT) models. Specifically, we implement three sets of AFT models, each set predicts the duration of different career stages: time to attain PI status, overall career length, and career length after becoming a PI. In each set, we fit an AFT model for each collaborative behaviour statistic, for both genders and the interaction between genders, just as we did when estimating the likelihood of becoming a PI in the previous section. The AFT model is written as follows:
@@
where $T_{i}$ denotes the time for a focal author to become a PI, their career length, or their career length after becoming a PI. The model includes $\beta_{0}$ as the intercept term, $\beta_{1}$ for the collaboration static effect, $\beta_{2}$ for the gender effect, and $\beta_{3}$ for the interaction effect between gender and the collaboration statistic. Finally, $\epsilon_{i}$ is an error term.
@@ 5.2.1 Overall career length
Figure 8 shows the interaction between gender and adjusted tie strength are found to be significant, suggesting that a weaker mean tie strength is associated with shorter careers, but this relationship is less pronounced for females compared to males. We see that overall having a larger network size appear to be associated with a longer academic career, while lower clustering coefficient and fewer coauthors per publication seem to be linked to a longer career length. These findings appear to be consistent between genders and indicate that a wider collaboration network, with weaker ties and clustering might contribute to career longevity.
Figure 8: The forest plot of adjusted collaboration statistics on career length. We present three types of model for each type of collaboration statistic (network size, tie strength, clustering coefficient and coauthors per publication). The interaction model refers to the full model described in Eq. (2) and includes the interaction effect of gender with the collaboration statistic. The female and male models assess the effect of the respective collaboration statistic on career length for the separate groups of female and male authors. The Number column shows the sample size of each model. The markers show the coefficient estimates, with error bars indicating the 95% confidence intervals (CIs) for each estimate.
@@ 5.2.2 Time to become a principal investigator and career length after PI
Figure 9 shows the interaction terms between gender and the four collaboration statistics are not statistically significant except between gender and clustering coefficient in career length after PI. All the adjusted collaboration statistics have a significant relationship with the time to becoming a PI (Figure 9). Specifically, we see that an increase in network size and number of coauthors per publication tends to prolong the expected time to becoming a PI. On the contrary, weaker mean tie strength and clustering coefficient seem to be associated with prolonged time to becoming a PI.
In Figure 9 we see that higher clustering between coauthors is associated with extended career in academia after attaining PI status, however the association is less pronounced for females. We see that larger network size and a greater number of co-authors per publication are positively correlated with career length post-PI, however this relationship is more ambiguous among the female authors due to the small sample size. Furthermore, a higher clustering coefficient may contribute to an extended career in academia after attaining PI status.
Figure 9: The forest plots of adjusted collaboration statistics on Time to become a PI and Career length after PI. In each forest plot, we present three types of model for each type of collaboration statistic (network size, tie strength, clustering coefficient and coauthors per publication). The interaction model refers to the full model described in Eq. (2) and includes the interaction effect of gender with the collaboration statistic. The female and male models assess the effect of the respective collaboration statistic on time to become a PI/career length after PI for the separate groups of female and male authors. The Number column shows the sample size of each model. The markers show the coefficient estimates, with error bars indicating the 95% confidence intervals (CIs) for each estimate.
@@ 6 Discussion
Here we have investigated the relationship between career progression and collaboration behaviour of male and female focal authors in physics. Our findings indicate that the relationship between collaboration behaviour and career progression is largely the same for both women and men. Specifically, the number of co-authors, the frequency at which authors publish with those co-authors, the degree to which their co-authors are interconnected, and the average size authors publish correlate with academic career progression in physics. We find that those who eventually reach the seniority of principal investigator (PI), tend to have published with more unique collaborators. In contrast, publishing repeatedly with the same highly interconnected collaborators and/or larger number of coauthors per publication is characteristic of shorter career lengths and those that do not attain PI status. These findings align well with the suggestions that prolonged engagement in specialised or internally-focused collaborations can become harmful [25]. However, at the same time, we also see that those that publish repeatedly with the same set of interconnected collaborators and do attain the status of PI, do so in less time. This is in agreement with recent findings on the benefits of highly repetitive collaborations for career progression [26].
We observe differences in collaboration behaviour and career progression between men and women. Women are more likely than men to publish repeatedly with the same coauthors and to participate in larger, more interconnected collaborations. These collaborative behaviours, regardless of gender, are linked to negative career progression outcomes. However, women exhibit these less favourable collaborative patterns more frequently and also have poorer career outcomes.
@@
Our results can potentially enhance existing policies that address gender balance in scientific collaboration. For instance, Horizon Europe, a major funding instrument in the European Union requests details regarding the “gender dimension” of all grant proposals [40]. First, policies could be implemented to facilitate opportunities for female researchers to engage with a wider array of collaborators outside their immediate circle, to align with our finding that a greater number of collaborators is positively associated with career progression. One major obstacle is the lower mobility of women compared to men, which substantially limits their network size [41]. Lowering restrictions on mobility might be achieved by providing hybrid working conditions, flexible multi-stage childcare support, or financial incentives that encourage mobility. Second, policies that reward impactful small-team publications could significantly enhance the visibility and recognition of female researchers’ contributions. These measures could include adjustments in performance evaluations to highlight individual work, dedicated funding for projects led by women, mentorship programs that support small teams, and workshops or training programs focused on developing skills for leading smaller research teams. Third, recognising and rewarding diverse collaboration efforts in performance evaluations and grant applications can motivate researchers to seek out varied and expansive professional relationships, ultimately supporting their career advancement.
@@
Here we have focused on statistics of the careers of authors in physics based only on publication data from the American Physical Society and therefore do not have a complete publication record for all these focal authors. While this necessarily means that we do not have the complete collaboration network for each author, it seems unlikely that collaboration behaviour would be significantly different when publishing in other journals. Additionally, inferring gender from names and principal investigator (PI) status from being the last author may be inaccurate due to factors such as non-binary names, cultural differences in naming conventions, and variations in authorship practices across fields and institutions. However, these methods are comparable to the state of the art and have been used in other studies [42, 43, 44, 9]. We find various associations between collaborative behaviours and career progression outcomes, however we cannot currently determine the direction of causality or even if a causal relation exists. Moreover, an author’s collaborative behaviour may not be consistent throughout different times of their career, which might not capture the intricacies and shifts in collaboration patterns across different career stages. For instance, an author might be highly collaborative during their early career and work more isolated later on, or vice versa. Such nuances get lost when looking at career-span averages. Future research could benefit from segmenting the career into different phases and analysing collaboration patterns within each phase to provide a more nuanced understanding.
Acknowledgements
@@
We obtain the dataset from the American Physical Society (APS) between 1893 to 2020. The dataset is comprised of all the publications in Physical Review, including 19 different journals: PRA, PRAB, PRAPPLIED, PRB, PRC, PRD, PRE, PRFLUIDS, PRI, PRL, PRMATERIALS, PRPER, PRRESERACH, PRSTAB, PRSTPER, PRX, PRXQUANTUM, and RMP.
Table S1: Summary of journals published by the American Physical Society (APS) from the years 1893 to 2020. The table shows the subject focus of each journal, its respective impact factor (IF) as per the 2022 Journal Citation Reports (Clarivate Analytics, 2022), and the total number of papers published in each journal up to 2020. The acronyms of the journals are provided alongside their full names for easy reference.
@@
PRA Physical Review A Atomic, Molecular, and Optical Physics 2.9 83,632
PRAB Physical Review Accelerators and Beams Accelerator science, Technology, and Applications 1.7 1,175
@@
PRD Physical Review D Particles, Fields, Gravitation, and Cosmology 5.0 94,833
@@
Name disambiguation is an important and often challenging first step in analysing scientists’ careers. The goal of name disambiguation is to accurately establish authorship links between authors and their respective publications using the available metadata. Here, we use an author’s name, affiliation and their collaborators to determine if two publications belong to the same or two distinct authors. We use a process based on prior research by Sinatra et al. who applied such an algorithm to the APS dataset [21]. The data is first filtered to exclude papers with more than $10$ authors and those without at least one affiliation per author in order to ensure sufficient contribution by each author. Subsequently, all author names are grouped if they have identical surnames and share matching initials. Sinatra et al. then merge the author identities of such author pairs if they meet one of the following criteria: (1) the authors cited each other at least once; (2) they shared at least one coauthor; (3) they shared at least one affiliation. To avoid ambiguities in the definition of the merging step, we follow a recently proposed implementation of this algorithm [10]. This solution follows the scheme provided by Sinatra et al. and merges matching author pairs only if their first names share the first initial and they are not conflicting with previously assigned first names.
@@
Since the APS dataset does not include author gender, we employ a recently developed method to infer genders based on authors’ names [22]. This method performs on-par with alternative solutions, but has the advantage that it follows open data and open source paradigms. That is, both source code and data are openly available. In addition, it was recently applied in a comparable study [10]. In an iterative procedure, the inference method assigns trust levels for their sources of data based on their agreement on gender assignments with the other sources. These trust scores then weight the decision on the final gender assignment of a given name. Repeating this procedure until both the gender inference decisions and the trust scores converge yields an estimate of the trustworthiness of a source, as well as a probabilistic gender assignment. We assign gender to author names when a certainty of at least $66\%$ can be guaranteed. Finally, we aggregate the assignment for multiple author names based on a majority vote. In case of a draw between male and female assignment, or when no gender could be inferred for any name, we assign an ‘unknown’ gender label. By following this rule, we identify the gender of $230,086$ authors.
@@
Initially, we construct separate models for each collaboration statistic, including network size, tie strength, number of links, clustering coefficient, and coauthors per publication. We use a linear regression model to predict network size, tie strength, and coauthors per publication because these collaboration statistics are continuous variables that can take on a wide range of values. We use a logistic regression model to predict the clustering coefficient since it ranges between 0 and 1. In these models, each collaboration statistic serves as the dependent variable, while the number of publications functions as the independent variable. This modelling allows us to isolate the impact of publication quantity on each aspect of collaboration. Subsequently, we extract the residuals from each model. These residuals represent the portion of collaboration statistics not explained by a linear relationship with the publication volume, thus effectively adjusting the collaboration statistics for the number of published papers. To facilitate comparability and eliminate scale discrepancies across different collaboration metrics, we standardise these residuals, ensuring they conform to a distribution with a mean of zero and a standard deviation of one.
@@
Examining the GLM coefficients also identifies some statistically significant differences in the mean collaboration statistics between genders. Table S2 indicates that female authors exhibit significantly higher adjusted tie strength ($\beta=0.142,p<0.01$), clustering coefficient ($\beta=0.101,p<0.01$), and coauthors per publication ($\beta=0.166,p<0.01$) compared to their male counterparts. These findings indicate that female authors tend to engage in more frequent and interconnected collaborations, as well as larger collaborative groups, than male authors.
@@
adj. network size adj. tie strength adj. clustering coefficient adj. coauthors per publication
@@
2 60259 37541 30130
3 51333 17763 8556
4 33288 6164 1387
5 19930 1884 167
6 12991 700 319
7 8615 336 2
8 6089 190 0.15
9 4070 97 0.01
To assess the impact of alphabetically ordered papers on the timing of becoming a PI, we divide PIs into two groups based on the first letter of their surname: Group A-M and Group N-Z. If Time to PI is not influenced by the initial letter of the surname; then ideally the distribution of Time to PI should be similar for both groups. We compare the distribution of Time to PI between these groups, both before and after removing alphabetically ordered papers. We find a discrepancy between Group A-M and Group N-Z in terms of Time to PI. Figure  S1 shows PIs with surnames towards the end of the alphabet appear to reach PI status earlier before removing alphabetically ordered papers, which could be attributed to them being listed last in alphabetically ordered papers. After removing alphabetically ordered papers (alphabetically ordered papers have at least 3 authors) from consideration, the distributions of Time to PI in both groups become similar.

### v2

@@ 5.2 Duration of career stages
We examine the impact of collaborative behaviour and gender on career progression using accelerated failure time (AFT) models. Specifically, we implement three sets of AFT models, each set predicts the duration of different career stages: time to attain PI status, overall career length and career length after becoming a PI. In each set, we fit an AFT model for each collaborative behaviour statistic, for both genders and the interaction between genders, just as we did when estimating the likelihood of becoming a PI in the previous section. The AFT model is written as follows:
@@
where $T_{i}$ denotes the time for a focal author to become a PI, their career length, or their career length after becoming a PI. The model includes $\beta_{0}$ as the intercept term, $\beta_{1}$ for the collaboration static effect, $\beta_{2}$ for the gender effect and $\beta_{3}$ for the interaction effect between gender and the collaboration statistic. Finally, $\epsilon_{i}$ is an error term.
@@ 5.2.1 Overall career length
Figure 8 shows the interaction between gender and adjusted tie strength is found to be significant, suggesting that a weaker mean tie strength is associated with longer careers, but this relationship is less pronounced for females compared to males. We see that overall having a larger network size appear to be associated with a longer academic career, while lower clustering coefficient and fewer co-authors per publication seem to be linked to a longer career length. These findings appear to be consistent between genders and indicate that a wider collaboration network, with weaker ties and clustering might contribute to career longevity.
Figure 8: The forest plot of adjusted collaboration statistics on career length. We present three types of model for each type of collaboration statistic (network size, tie strength, clustering coefficient and co-authors per publication). The interaction model refers to the full model described in Eq. (2) and includes the interaction effect of gender with the collaboration statistic. The female and male models assess the effect of the respective collaboration statistic on career length for the separate groups of female and male authors. The Number column shows the sample size of each model. The markers show the coefficient estimates, with error bars indicating the 95% confidence intervals (CIs) for each estimate.
@@ 5.2.2 Time to become a principal investigator and career length after PI
Figure 9 shows the interaction terms between gender and the four collaboration statistics are not statistically significant except between gender and clustering coefficient in career length after PI. The adjusted collaboration statistics have a significant relationship with the time to becoming a PI (Figure 9). Specifically, we see that an increase in network size and number of co-authors per publication tends to prolong the expected time to becoming a PI. On the contrary, weaker mean tie strength and clustering coefficient seem to be associated with prolonged time to becoming a PI for male authors.
In Figure 9 we see that a larger network size, a higher clustering coefficient and a greater number of co-authors per publication are positively correlated with career length post-PI, but this relationship is more ambiguous among the female authors due to the small sample size. However, our additional analyses (given in supplementary material S5) suggest that having a larger network size, higher clustering coefficient and greater average number of co-authors at the time of becoming a PI is negatively associated with career length post-PI. This suggests that tightly-knit or larger-scale collaborations at the time of becoming PI might limit opportunities for sustained academic development, even though final higher clustering may contribute to extending a career in some contexts. In addition to these collaboration statistics, we also explored the gender composition of ego networks as a supplementary collaboration statistic (see more details in Appendix  S8). We find that female authors tend to collaborate with a higher proportion of female co-authors and that female ratio in co-authorship networks shows a weak association with career outcomes.
Figure 9: The forest plots of adjusted collaboration statistics on Time to become a PI and Career length after PI. In each forest plot, we present three types of model for each type of collaboration statistic (network size, tie strength, clustering coefficient and co-authors per publication). The interaction model refers to the full model described in Eq. (2) and includes the interaction effect of gender with the collaboration statistic. The female and male models assess the effect of the respective collaboration statistic on time to become a PI/career length after PI for the separate groups of female and male authors. The Number column shows the sample size of each model. The markers show the coefficient estimates, with error bars indicating the 95% confidence intervals (CIs) for each estimate.
@@ 6 Discussion
Here we have examined the collaboration behaviour and career progression of female and male focal authors in physics. Our findings indicate that the relationship between collaboration behaviour and career progression is largely the same for both women and men. However, despite this relationship being consistent across genders, we do observe differences in the way men and women collaborate and how they progress in their careers. Women are more likely than men to publish repeatedly with the same co-authors and to participate in larger, more interconnected collaborations, a pattern also observed in economics [26]. These collaborative behaviours, regardless of gender, are linked to negative career progression outcomes. However, women exhibit these less favourable collaborative patterns more frequently and also have poorer career outcomes.
We find that those who eventually reach the seniority of principal investigator (PI), tend to have published with more unique collaborators. In contrast, publishing repeatedly with the same highly interconnected collaborators and/or larger number of co-authors per publication is characteristic of shorter career lengths and those that do not attain PI status. These findings align well with the suggestions that prolonged engagement in specialised or internally-focused collaborations can become harmful [27] and that publishing with a large number of co-authors per publication may dilute individual contributions and visibility [28]. However, at the same time, we also see that those that publish repeatedly with the same set of interconnected collaborators and do attain the status of PI, do so in less time. This is in agreement with recent findings on the benefits of highly repetitive collaborations for career progression [29].
Our findings align with prior research, such as the study by Van der Wal et al. [9], in highlighting the importance of collaboration behaviour in relation to career progression. Lindenlaub and Prummer [30] demonstrate that tighter networks hinder productivity by restricting information flow and novel idea exchange, factors critical for career progression. Both studies observe that larger and more diverse networks are beneficial for career advancement, while tightly clustered networks (indicated by higher clustering coefficients) may limit diversity in collaborations and visibility within broader professional communities. Our observation that smaller, tightly interconnected networks are associated with reduced likelihood of transitioning to PI status is consistent with these studies. Moreover, this pattern remains consistent across subfields in physics (see S6 for further details). However, we also identify that more clustered networks have an accelerated transition to PI status in physics. In contrast, the study by Van der Wal et al. associates less clustered networks with faster PI transitions in behavioural ecology. This difference may reflect variations in the collaborative and publishing norms of these disciplines. Furthermore, while Van der Wal et al. identify stronger positive relationships between network statistics and career progression for women, our results do not indicate a significant difference between genders.
@@
Our results can potentially enhance existing policies that address gender balance in scientific collaboration. For instance, Horizon Europe, a major funding instrument in the European Union requests details regarding the “gender dimension” of all grant proposals [44]. First, policies could be implemented to facilitate opportunities for female researchers to engage with a wider array of collaborators outside their immediate circle, to align with our finding that a greater number of collaborators is positively associated with career progression. One major obstacle is the lower mobility of women compared to men, which substantially limits their network size [45]. Lowering restrictions on mobility might be achieved by providing hybrid working conditions, flexible multi-stage childcare support, or financial incentives that encourage mobility. Second, policies that reward impactful small-team publications could significantly enhance the visibility and recognition of female researchers’ contributions. These measures could include adjustments in performance evaluations to highlight individual work, dedicated funding for projects led by women, mentorship programs that support small teams and workshops or training programs focused on developing skills for leading smaller research teams. Third, recognising and rewarding diverse collaboration efforts in performance evaluations and grant applications can motivate researchers to seek out varied and expansive professional relationships, ultimately supporting their career advancement.
@@
Here we have focused on statistics of the careers of authors in physics based only on publication data from the American Physical Society and therefore do not have a complete publication record for all these focal authors. While this necessarily means that we do not have the complete collaboration network for each author, it seems unlikely that collaboration behaviour would be significantly different when publishing in other journals. Additionally, inferring gender from names and principal investigator (PI) status from being the last author may be inaccurate due to factors such as non-binary names, cultural differences in naming conventions and variations in authorship practices across fields and institutions. However, these methods are comparable to the state of the art and have been used in other studies [46, 47, 48, 9]. We find various associations between collaborative behaviours and career progression outcomes, however we cannot currently determine the direction of causality or even if a causal relation exists. Moreover, an author’s collaborative behaviour may not be consistent throughout different times of their career, which might not capture the intricacies and shifts in collaboration patterns across different career stages. For instance, an author might be highly collaborative during their early career and work more isolated later on, or vice versa. Such nuances get lost when looking at career-span averages. Future research could benefit from segmenting the career into different phases and analysing collaboration patterns within each phase to provide a more nuanced understanding.
Acknowledgment
@@
Data availability
The APS datasets are available upon request to the American Physical Society (https://journals.aps.org/datasets). The data analysed in this study have been deposited in the Dryad Digital Repository and can be accessed via https://doi.org/10.5061/dryad.b5mkkwhr4. Relevant code for this research work are stored in GitHub: https://github.com/mingrongshe/Gender_differences_Physics and have been archived within the Zenodo repository: https://doi.org/10.5281/zenodo.13739434. Before publication, the code will be further developed to generate and integrate these intermediate results.
@@
We obtain the dataset from the American Physical Society (APS) between 1893 to 2020. The dataset is comprised of all the publications in Physical Review, including 19 different journals: PRA, PRAB, PRAPPLIED, PRB, PRC, PRD, PRE, PRFLUIDS, PRI, PRL, PRMATERIALS, PRPER, PRRESERACH, PRSTAB, PRSTPER, PRX, PRXQUANTUM and RMP.
Table S1: Summary of journals published by the American Physical Society (APS) from the years 1893 to 2020. The table shows the subject focus of each journal, its respective impact factor (IF) as per the 2022 Journal Citation Reports (Clarivate Analytics, 2022) and the total number of papers published in each journal up to 2020. The acronyms of the journals are provided alongside their full names for easy reference.
@@
PRA Physical Review A Atomic, Molecular and Optical Physics 2.9 83,632
PRAB Physical Review Accelerators and Beams Accelerator science, Technology and Applications 1.7 1,175
@@
PRD Physical Review D Particles, Fields, Gravitation and Cosmology 5.0 94,833
@@
Name disambiguation is an important and often challenging first step in analysing scientists’ careers. The difficulty arises from the fact that multiple individuals can share the same name, while a single individual may appear under different name variants. The goal of name disambiguation is to accurately establish authorship links between authors and their respective publications using the available metadata. Here, we use an author’s name, affiliation and their collaborators to determine if two publications belong to the same or to different identities. We use a process based on prior research by Sinatra et al. who applied such an algorithm to the APS dataset [22]. The data is first filtered to exclude papers with more than $10$ authors and those without at least one affiliation per author in order to ensure sufficient contribution by each author. Subsequently, all author names are grouped if they have identical surnames and share matching initials. Sinatra et al. then merge the author identities of such author pairs if they meet one of the following criteria: (1) the authors cited each other at least once; (2) they shared at least one co-author; (3) they shared at least one affiliation. To avoid ambiguities in the definition of the merging step, we follow a recently proposed implementation of this algorithm [10]. This solution follows the scheme provided by Sinatra et al. and merges matching author pairs only if their first names share the first initial and they are not conflicting with previously assigned first names.
@@
Since the APS dataset does not include author gender, we employ a recently developed method to infer genders based on authors’ names [23]. This method performs on-par with alternative solutions, but has the advantage that it follows open data and open source paradigms. That is, both source code and data are openly available. In addition, it was recently applied in a comparable study [10]. In an iterative procedure, the inference method assigns trust levels for their sources of data based on their agreement on gender assignments with the other sources. These trust scores then weight the decision on the final gender assignment of a given name. Repeating this procedure until both the gender inference decisions and the trust scores converge yields an estimate of the trustworthiness of a source, as well as a probabilistic gender assignment. We assign gender to author names when a certainty of at least $66\%$ can be guaranteed. Finally, we aggregate the assignment for multiple author names based on a majority vote. In case of a draw between male and female assignment, or when no gender could be inferred for any name, we assign an ‘unknown’ gender label. By following this rule, we identify the gender of $248,788$ authors.
To ensure the robustness of our findings, we also reanalysed the data using a stricter $95\%$ probability threshold for gender assignment. Despite the reduced sample size, the main findings and conclusions remain consistent with those obtained using the $66\%$ threshold. Additionally, we performed a separate analysis using gender data derived from Genderize and Gender API. The results from this independent method were largely consistent, with one exception: the interaction term between gender and tie strength was not significant(see Figure S1).
Figure S1: The forest plots of adjusted collaboration statistics on likelihood of being a PI and career length using different gender inference methods. The plots compare results obtained under two gender inference methods. (a) shows the analysis with a stricter 95% probability threshold for gender assignment, while (b) shows results derived from Genderize and Gender API. In each forest plot, we present three types of model for each type of collaboration statistic (network size, tie strength, clustering coefficient and co-authors per publication). The interaction model refers to the interaction effect of gender with the collaboration statistic. The female and male models assess the effect of the respective collaboration statistic on likelihood of being a PI/career length for the separate groups of female and male authors. The Number column shows the sample size of each model. The markers show the coefficient estimates, with error bars indicating the 95% confidence intervals (CIs) for each estimate.
@@
Initially, we construct separate models for each collaboration statistic, including network size, tie strength, clustering coefficient and co-authors per publication. We use a linear regression model to predict network size, tie strength and co-authors per publication because these collaboration statistics are continuous variables that can take on a wide range of values. We use a logistic regression model to predict the clustering coefficient since it ranges between 0 and 1. In these models, each collaboration statistic serves as the dependent variable, while the number of publications functions as the independent variable. This modelling allows us to isolate the impact of publication quantity on each aspect of collaboration. Subsequently, we extract the residuals from each model. These residuals represent the portion of collaboration statistics not explained by a linear relationship with the publication volume, thus effectively adjusting the collaboration statistics for the number of published papers. To facilitate comparability and eliminate scale discrepancies across different collaboration metrics, we standardise these residuals, ensuring they conform to a distribution with a mean of zero and a standard deviation of one.
@@
Examining the GLM coefficients also identifies some statistically significant differences in the mean collaboration statistics between genders. Table S2 indicates that female authors exhibit significantly higher adjusted tie strength ($\beta=0.173,p<0.01$), clustering coefficient ($\beta=0.154,p<0.01$) and co-authors per publication ($\beta=0.248,p<0.01$) compared to their male counterparts. These findings indicate that female authors tend to engage in more frequent and interconnected collaborations, as well as larger collaborative groups, than male authors.
@@
adj. network size adj. tie strength adj. clustering coefficient adj. co-authors per publication
@@
2 52467 32836 26234
3 44626 15588 7438
4 28809 5392 1201
5 17351 1690 145
6 11376 630 16
7 7502 316 1.49
8 5338 180 0.13
9 3553 95 0.01
To assess the impact of alphabetically ordered papers on the timing of becoming a PI, we divide PIs into two groups based on the first letter of their surname: Group A-M and Group N-Z. If Time to PI is not influenced by the initial letter of the surname; then ideally the distribution of Time to PI should be similar for both groups. We conducted Wilcoxon signed-rank tests to assess whether the median Time to PI differed significantly between the two alphabetical groups (A-M and N-Z), both before and after accounting for alphabetically ordered papers. The results showed a significant median difference between the two groups before adjustment (p = 0.0097) but not after adjustment (p = 0.1792). We also compare the distribution of Time to PI between these groups, both before and after removing alphabetically ordered papers. We find a discrepancy between Group A-M and Group N-Z in terms of Time to PI. Figure S2 shows PIs with surnames towards the end of the alphabet appear to reach PI status earlier before removing alphabetically ordered papers, which could be attributed to them being listed last in alphabetically ordered papers. After removing alphabetically ordered papers (alphabetically ordered papers have at least 3 authors) from consideration, the distributions of Time to PI in both groups become similar.